Minutes of February 10, 2011 Meeting
- Present:
- G.Sandberg, B. Lipinski, D. Meister, J.Gardner, T.Meyer, M.Tokar, P. Weniger, J. Teichmann, A.Martinez, M.Harris-Clark, E. Fizur, E. Suarez, R. Barbalace, K. Haviland,V. Cvetkovic, J. Reinhart, D. Voorhees, Y. Ishii
1. Update on the new Relais EZB
The Relais Implementation Team conducted one training session in New Brunswick for about 10 staff,
covering how to print book bands, ship (fill) requests, set requests to unfilled, and query
requests. We are still unable to Receive, Return, and Complete requests. Until we can test these
functions, training for the IDS/UPS site libraries will not be re-scheduled. IIS is working on new
NetID and password login to EZB. ILS staff is looking at ways to continue to convert ILLs to EZBs
on the patron's behalf. PALCI 's current go live date is the week of Feb. 14th but it will
undoubtedly be postponed again.
2. Revised EZB webpage
The group made suggestions for the webpage. Judy will incorporate additional changes and post
another revised version on sakai.
3. Proposed new UMDNJ courtesy notices
The notices are similar to what we send to Emeritus and Community borrowers when their annual
privileges are about to expire. They are useful reminder. 8 notices have been created for
faculty/staff and students with or without remote access and with or without checkts.
4. Plastic Guest Borrower cards
It's time to order a new supply of plastic Guest borrower cards. The group reviewed the current
design and made suggestions for name, status, and barcode labels, and recommended we use one
plastic card for guests and occasional primary borrowers. Judy will draft a re-design and ask for
further input.
5. Information Sharing
- Vib thanked Michele for traveling to Camden and the great training she provided new staff there.
- Dean reported the Annex has experienced roof leaks in the stacks, a roofing company came and did repairs. Facilities have cleared a blocked Emergency exit in the stacks; lost ½ of student staff and hired a voucher back to fill in; 150 boxes ready for Recyclomania. Moving Alex microfilm boxes to better shelves for permanent storage. Transfers are mostly from East Asian at the moment. RUPD need to re-do a security walkthrough of the Annex for the personnel security audit.
- Glenn reported that Mary Belasco is retiring next month.
- Tracey relayed an update from Bob Warwick-the patron load on 1/11 had no FACSTF records loaded but they resumed loading on 1/15. All issues regarding the new source of patron records have been resolved. Zimbra migration is at the end phase. If anyone has not yet been migrated, please contact Tracey.
- Maggie asked about a contact for barcodes and Tracey asked her to send an email message to bar_codes@email.rutgers.edu
- The Reserves group is gathering statistics. Physics and Chemistry have completed shifts. SERC's 3-M exit gate has been replaced with a people counter.
- Douglass has been busy with users and events. Media Center and Fordham Lab had an open house on 2/1/11 that was well attended by over 40 faculty and staff.
- Student Coordinators have been busy trying to adjust to the new RIAS payroll systems. Students have adjusted to the new WebClock with few problems. Rutgers Day planning is fast approaching.
- Kilmer is in the process of reintroducing a Rec Reading collection.
- RSG group is waiting patiently for the new EZB system.
- The former serials counter and old imaging services office counter were removed from the 3rd floor of LSM.
- A new reference desk was installed at Alexander and public and RU computer moved. Café construction may start in March; the current periodicals stacks have been removed and moved to the old wing, tables are moving to Art and LSM.
- Kristen Ko is back at Art. Art library sprinkler burst a pipe but no books got wet.
- The Training Group is working on producing short training videos on a variety of topics. Also, reviewing the borrower's brochure and looking at a new format. The credit card format has been eliminated since we no longer issues PINS to primary users.
- Eddie will be evaluating the current periodical shelving at Alex to determine what we might keep or send to surplus.
- Imaging Services has begun to plan for the next round of equipment leases which will include a larger number of public scanners. Administrative multifunction copiers will also be included. Chang's copier, print release station will be moved today to fit with plans to reassign space there.
- Construction is still causing some confusion and disruption at Robeson, the OIT lab will be closed from Friday afternoon until Wed. Eric is busy moving resources and equipment around; snow meliting causing new roof leaks.
- Questions were asked about updating standard RDS replies and canned AAL responses, with regard to the change in PIN policy and NetIDs.
- Comments about meeting on sakai chat: liked it; like in-person better, or teleconference; hard to keep track of conversations and responses; OK once we got going, miss face-to-face; fine for brief agenda topics; good for lighter agenda meetings; good because of limited parking due to snow; flow and give & take difficult in chat; OK, just not all the time; would be nice to do some screen sharing; like it for a lighter agenda; different, the pace, flow changed throughout but overall enjoyable; good for chatting but not for making decisions in a large group; multiple threads going at the same time are difficult to track.